A liquid is kept in a  vessel of cylindrical size which is rotating along its axis. The liquid rises at the sides. If the radius of the vessel is 0.05m and the speed of rotaion is 2 rev/second , find the difference in height of the liquid at the centre of the vessel and its side g=9.8 m/s2

Grade:11

1 Answers

Arun
6 years ago
Please find below the solution to your asked query.

according to the bernoulli's theorem:p+12ρv2 = constantwe knoe v = rωapplying these equation at the side wall and centre of the vesselpcentre 12ρvcentre2 =  pside 12ρvside2 vcenter = 0pcentre pside = 12ρr2ω2hρg =12ρr2ω2h = 12gr2ω2  = 1×0.05×0.05×(2π×2)220 = 0.02 m= 2 cm